Two years ago a man was 7 times as old as his son, but in three years he will only be four times as old as the boy. How old is each now?
Answer by richwmiller(17219)

m-2=7*(s-2)
m+3=4(s+3)
m=37 s=7
check
35=7*5
40=4*10
ok